OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (a1234):

x^2 – 36 = 0 and x^2 = 8x – 12 x^2 = 36 Sqrt 36 = -6 or 6------> roots. x^2 = 8x – 12 x^2 - 8x + 12 = 0 Find two numbers that have a product of 12 and a sum of -8. x^2 - 6x - 2x + 12 = 0 x(x - 6) - 2(x - 6) = 0 (x - 2)(x - 6) = 0 x = 2 x = 6 ----> roots